# Cross-domain Multi-modal Few-shot Object Detection via Rich Text

### v1

@@ Abstract
Cross-modal feature extraction and integration have led to steady performance improvements in few-shot learning tasks due to generating richer features. However, existing multi-modal object detection (MM-OD) methods degrade when facing significant domain-shift and are sample insufficient. We hypothesize that rich text information could more effectively help the model to build a knowledge relationship between the vision instance and its language description and can help mitigate domain shift. Specifically, we study the Cross-Domain few-shot generalization of MM-OD (CDMM-FSOD) and propose a meta-learning based multi-modal few-shot object detection method that utilizes rich text semantic information as an auxiliary modality to achieve domain adaptation in the context of FSOD. Our proposed network contains (i) a multi-modal feature aggregation module that aligns the vision and language support feature embeddings and (ii) a rich text semantic rectify module that utilizes bidirectional text feature generation to reinforce multi-modal feature alignment and thus to enhance the model’s language understanding capability. We evaluate our model on common standard cross-domain object detection datasets and demonstrate that our approach considerably outperforms existing FSOD methods. Our implementation is publicly available: https://github.com/zshanggu/CDMM
Keywords:
Few-shot Object Detection Multi-modal Cross-domain
@@ 1 Introduction
The goal in few-shot object detection (FSOD) is to train a model that can generalize to detecting objects for which we only have a few labeled examples [56, 46]. This task has applications in domains where encountering rare or less common objects is frequent, such as autonomous driving. A typical FSOD training method follows a two-fold training procedure. First, a model is pre-trained on a number of base classes for which we have abundant data and then the model is fine-tuned on a set of novel classes with few annotated samples. In the classic setting, the base and the novel classes are generally within the same feature domain. Existing FSOD methods primarily rely on either fine-tuning or meta-learning paradigms [56]. Fine-tuning usually requires less GPU resources but have less generalization power. In contrast, methods based on meta-learning have more robust generalization because of the class-agnostic training process and usually make more accurate inference but demand more GPU resources for building task episodes [60]. Multi-modal few-shot object detection (MM-FSOD), as a new approach for FSOD, aims at utilizing extra text information to enhance visual feature representation and provide potential zero-shot inference ability (see Fig. 1). Broadly speaking, multi-modal machine learning involves training on diverse sources of information, which can range from vision, language, acoustic, tactile, and others. Multi-modal learning has shown significant success in problems such as visual question-answering [10, 49], robotics [3, 30], continual learning  [4], and medical image analysis [27]. Multi-modality is also central to modern Generative AI models such as GPT-4 [37] and Gemini [14] which can process and produce text and images to perform complex tasks.
@@
Figure 1: Different FSOD tasks. The classic FSOD task (top) involves only visual information. MM-FSOD (middle) introduces a language modality, which can provide extra information to improve FSOD performance. Our proposed CDMM-FSOD task (bottom) is tailored for cross-domain scenarios and extends MM-FSOD to use richer text and semantics. Above, we show a cross-domain example where a model might train on images and text of common data (such as birds) and would need to generalize to detection of less common, substantially different data (such as electrical circuits).
Despite the success of multi-modal object detection (MM-OD) methods on few-shot task, the source and the target data may have a substantial domain gap in some practical detection problems. Thus, MM-OD methods may have degraded performance in such cross-domain, data scarce scenarios. In Fig. 2, we visualize our problem description, where we train and test existing object detection methods on data where the source and target domains differ. We propose the cross-domain multi-modal few-shot object detection (CDMM-FSOD) task, aiming to bridge the domain gap while transferring information from the source domain to the target domain. We hypothesize that using rich text within a sentence can be more effective than relying solely on simple semantics. This property may be particularly true when images in the target domain require technical language to describe, since most models are trained on data common in daily life. For example, when detecting product defects or analyzing medical images, experts could use a professional training manual to learn how to distinguish the appearance of the objects and how to label the image. However, for neural network models, how to utilize such rich text information in helping with these detection task is still unclear. We aim to investigate whether and how rich text can help address such cross-domain data gaps. In Fig. 1, we illustrate the difference of our proposed CDMM-FSOD compared to classic FSOD and MM-FSOD. Our rich text contain detailed and technical descriptions of the training data, and our novel data (the electrical circuit in this example) has prominent domain shift.
@@
Figure 2: Performance results on 10-shot object detection on multiple cross-domain, few-shot datasets. We observe substantial cross-domain degradation for existing multi-modal object detection methods Next-chat [59], GOAT [55], and ViLD [15], as well as for a single-modal detection method, Meta-DETR [60]. In contrast, our proposed method has stronger performance on out-of-domain data.
@@
We choose a meta-learning version of Detection Transformer (DETR) [5], i.e. Meta-DETR [60] as our baseline because of its Transformer-based structure and excellent modality adaption capacity. Unlike the classic multi-modal setting, we define a rich text description for each training category which may contain technical terminology. These sentences function as the linguistic support set for our meta-learning process, analogous to the role played by the image support set in terms of functionality. We design a meta-learning multi-modal aggregated feature module to fuse the vision and the language embedding and map them to a class-agnostic feature space as the meta support feature. Furthermore, to ensure the model obtains the detailed knowledge from the rich semantics, we propose a rich semantic rectify module to align the generated language embedding with the ground truth language embedding. In our experiments, we demonstrate consider performance improvement by using rich text. We further discuss and investigate the effect of text length on performance.
Our contributions include:
  • To the best of our knowledge, we are the first to propose and study the cross-domain problem for MM-FSOD.
@@
  • We propose a rich semantic rectify module that learns to align the generated language embedding and ground truth language embedding, enabling the model to better understand the text.
  • Our experiments on cross-domain detection data demonstrate that our approach outperforms existing methods, suggesting that multi-modal learning, when combined with rich text, is effective for detection of few-shot, out-of-domain data.
@@ 2.1 Few-shot Object Detection
The FSOD task is to train a model to accurately detect novel categories where only a few labeled samples are available. The dominant training paradigms of FSOD include methods based on fine-tuning and meta-learning. Fine-tuning methods often first pre-train a model on base data with sufficient training samples for each category, and then directly fine-tune the model on novel categories with few samples. Representative works include TFA [56], FSCE [50], FSRC [45], DeFRCN [38], etc. In meta-learning methods, i.e. learning to learn, instead of aiming at specific classes, a model is trained to learn a metric ability through tasks (i.e. episodes), so that it can utilize a Siamese network [28] to achieve accurate classification by measuring the similarity between support and query features; thus, it is expected to learn class-agnostic knowledge [6]. Each training episode in meta-learning methods builds a support set and query set; an $n$-way $k$-shot support set consists of $n$ categories with $k$ training instances for each. Such work includes Meta-FRCN [18] and Meta-DETR [60].
To avoid catastrophic forgetting of the base categories, researchers tune FSOD models on both base and novel data. However, while fine-tuning methods directly transfer the model on novel data in a class-specific manner, meta-learning methods tend to result in class-agnostic models, because they aim at learning a metric scheme between query features and support features. TFA [56] demonstrates that fine-tuning methods have potential to surpass meta-learning methods in detection performance. However, FCT [20] shows that meta-learning is still a powerful solution for FSOD, and that such methods are better at rapidly adapting to new tasks, due to their task-level feature representation.
The traditional backbone networks for FSOD include Faster R-CNN [42], YOLO [41], and ViT [8]. More recently, Detection Transformer (DETR) [5], a novel object detection network, is gaining increasing popularity for FSOD due to its competitive performance compared to classical frameworks; furthermore, its transformer-based architecture facilitates integration with NLP tasks. In our work, we thus follow the meta-learning paradigm provided by Meta-DETR [60].
@@ 3.2 Rich Text
We propose to manually define a fixed, rich text description for each category $c\in C_{B}\cup C_{N}$ to form the text $\ell$ for the training data. We denote the set of text descriptions as $W^{C}=\{w^{1},w^{2},\ldots,w^{C}\}$, where $C=|C_{B}\cup C_{N}|$ is the number of categories. Each text description contains multiple tokens with variable length of $M$, noted as $w^{i}=\{w^{i}_{1},w^{i}_{2},\ldots,w^{i}_{M}\}$, where the first token $w^{i}_{1}=\langle s\rangle$ is the start of the sentence, and $w^{i}_{M}=\langle/s\rangle$ is the end of the sentence. To form the set $W^{C}$, we use two strategies: (i) we refer to the Wikipedia content of each category and define each category from the aspects of color, shape, attribute, material, etc. (ii) besides the aspects from the first strategy, we extend the sentence by adding the depiction of the image context while keep other text unchanged. For example, the original sentence “Motorbikes have two wheels, a motor, and a sturdy frame” will be extended as “Motorbikes have two wheels, a motor, and a sturdy frame, usually ride by a person.” In the revised sentence, we introduce a common visual relationship between two categories: motorbikes and persons. Please refer to the Appendix for our complete text descriptions for all classes.
Figure 3: The overall structure of our model. Our proposed multi-modal feature aggregation module and rich text rectify module are in red blocks, the detailed structures of them are in Fig. 4. The multi-modal feature aggregation module is response for the cross-modality feature embedding mix. The rich text rectify module is used to reinforce the cross-modality understanding of the model. We design this end-to-end model that takes a set of support and query images, as well as a group of rich category text as input for training, and outputs the object detection results of the query images.
@@ 3.3 Proposed Architecture
We use Meta-DETR [60] as our baseline network, which is built upon deformable DETR [64]. DETR is a transformer-based [53] object detection model. It has a ResNet [22] backbone to obtain the high-level image features; and then an encoder is applied to get the encoder memory embedding; followed by a decoder that could convert the memory embedding back to the object proposal. For the backbone part, we follow Meta-DETR [60] and use a weight-shared self attention module to build the meta-learning scheme for the query and support images. The classic DETR network includes an encoder and a decoder module, both consists of six multi-head attention layers. Meta-DETR build the meta-learning module only at the first multi-head attention layer of the encoder module. For the meat-learning module, the query and support images will be firstly sent to a shared ResNet feature extractor to get the basic image features. Then the query and support features will be passed through a multi-head attention module so that they are in the same feature space. The support sample feature will then serve as the $K$ and $V$ matrices while the query image feature is $Q$ of a single-head self-attention module. This single-head self-attention module is response for the class-specific feature matching among the support and query feature. Meanwhile, in order to align with the class-agnostic learning task setting of meta-learning, this single-head self-attention module will be applied on a parallel branch to map the class-specific ‘support’-‘query’ relationship to a class-agnostic ‘task embedding’-‘query’ relationship. Superficially, Meta-DETR designs an independent trainable class-agnostic feature prototype as $V^{\prime}$ for the parallel single-head self-attention branch. And the support and query sample feature will still serve as $K^{\prime}$ and $Q^{\prime}$ respectively. In this way, the single-head self-attention module will achieve an important feature mapping so that the original class-specific support sample feature could be converted to class-agnostic prototype feature. During the test time, the model will only use the class-agnostic prototype as support features. And then the aggregated class-agnostic prototype will be sent to the input of DETR to generate the final detection output.
@@
Figure 4: The upper block is the detail of our meta-learning multi-modal aggregation module and the lower on is the rich semantic rectify module. We use different colors for different feature branches. Please note that the rich semantic rectify module is for training process only and thus will not participate in the inferencing process.
@@
The overall architecture of our proposed method is shown in Fig. 3. We add the rich text label as the auxiliary support feature. Then we aggregate the image and language support feature to train the single-head self-attention meta-learning module. Here we note this module as meta-learning multi-modal feature aggregation module. Besides, in order to ensure that the model properly learn the knowledge from the rich text information, we adopt a rich text semantic rectify module to bidirectionally generate language feature, and then enforce it to align with the ground-truth language feature.
3.4 Multi-modal Feature Aggregation
Multi-modal feature aggregation (MM Aggre.) is the core module to fuse the multi-modal information, which is capable of aggregating multiple support categories. The detailed architecture of this module is visualized in Fig. 4. we use the existing CLIP pre-trained model to extract the primary image feature of the support/query set and the primary semantic feature of the rich text. Then the support image feature will be processed by a RoIAlign layer and an average pooling layer to get the instance-level category prototypes. This is a common operation in object detection because the image feature could not represent the instance feature; there usually exists multiple foreground objects in one image. Simultaneously, similar to prior object detection works, the background, as an extra hidden label, is also initialized as one of the category prototypes. And then these features will be mapped to the same feature space through a multi-head self-attention module (i.e. the first DETR encoder layer).
$A=\text{softmax}\left(\frac{QS^{T}}{\sqrt{d}}\right)$ (1)
$Q_{1}=A\sigma(S)\odot Q,$ (2)
where $\odot$ represents element-wise product. The output $Q_{1}$ is a refined query feature map that keep only the ‘foreground’ feature.
For the ‘task embedding’-‘query’ encoding matching, a set of class-agnostic task prototype ($T\in\mathbb{R}^{(C+1)\times d}$) is defined to replace the support class prototype. In other words, the class-specific support class prototype will be mapped to the pre-defined task prototype, so that the query feature will be matched with the task prototypes for the prediction process. The task prototype $T$ would serve as the $V^{\prime}$ matrix for the same single head self-attention module as the ‘support’-‘query’ feature mapping. The initialization of $T$ is similar to sinusoidal position embedding in transformer. The corresponding output will be:
$Q_{2}=AT$ (3)
3.5 Rich Text Semantic Rectify
The rich text semantic rectify module (Rich Text Rect.) is depicted in Fig. 4. Our idea is to use a transformer encoder structure  [1], noted by $f_{\theta}$ to generate the category-level language sequence features of the fused support-query samples, and then the generated language feature is expected to be aligned with its ground truth feature. Additionally, this language generation process is bidirectional, which means we generate the sentence in both forward and backward direction, so that we ensure the model have robust image-text understanding ability.
For a given category $c$, its rich text tokens undergo refinement through the multi-head self-attention module that shared with the vision branch to get its semantic prototype $l^{c}$, as shown in Fig. 4. Additionally, the support and query features are averaged to create a composite feature $p$, which serves as the $K$ and $V$ matrices for a multi-head self-attention module $A_{3}$, as shown in Fig. 4. In $A_{3}$, the refined text tokens act as the $Q$ matrix. Subsequently, it generates the predicted text sequence $\hat{W}$ of length $M$. This entire process is duplicated for both forward and backward prediction components, i.e. the model will predict $\hat{W}$ from both left to right and right to left direction so that it could understand the rich text through its bidirectional context. The loss function is:
$\mathcal{L}_{rect}=\frac{1}{2}\left[\sum_{i=1}^{C}\sum_{j=2}^{M}-\mathrm{log}f_{\theta}(\hat{w}_{ij}-w_{ij}|p)+\sum_{i=1}^{C}\sum_{j=1}^{M-1}-\mathrm{log}f_{\theta}(\hat{w}_{ij}-w_{ij}|p)\right]$ (4)
4 Experiments
4.1 Implementation Details
We conduct our experiments on hardware consisting of 8 3090 GPUs for parallel training and testing. We borrow hyper-parameters from DETR [5] by setting the initial learning rate as 0.001, and using a batch size of 4 for 1-shot scenarios and 1 for 5-shot and 10-shot scenarios. The DETR encoder and decoder both have 6 self-attention layers (the first layer of the encoder is embedded in our meta-learning multi-modal aggregation module). We follow Afham et al. [1] for the architecture of the image-text transformer decoder.
[…]
@@ 4.1 Implementation Details
4.2 Datasets
Following the standard evaluation on the CD-FSOD benchmark provided by Xiong et al. [57], we use ArTaxOr [9], DIOR [33] and UODD [25] as our target domain few-shot datasets. ArTaxOr contains 7 arthropods categories, DIOR consists of 20 satellite images categories, and UODD has 3 underwater creature categories. We also provide results for the CD-FSOD benchmark proposed by Lee et al. [29] in the Appendix.
Furthermore, we test on three other recently introduced CD-FSOD benchmarks that adopt balanced sampling: Clipart1k [24], DeepFish [44], and NEU-DET [48]. Since their corresponding data splits for CD-FSOD are not officially published, we put the results in the Appendix.
To measure the generalization of our method, we evaluate on the general FSOD benchmark of PASCAL VOC. Following the pioneering evaluation pipeline for FSOD [56], 15 of the categories of PASCAL VOC are selected as base classes for pre-training, and the remaining 5 categories are used as the novel classes for few-shot fine-tuning. See the Appendix for a detailed description.
4.3 Results on CD-FSOD Datasets
As introduced in Sec. 3.2, we build rich text manually or by using an LLM [52]. As indicated in Tab. 1, either way could generate significant improvement on the out-of-domain benchmarks. It is observed that LLM-generated rich text brings better performance on these three benchmarks, because the generated text contains more comprehensive sentences. However, this is not a universal situation, because sometimes the LLM failed to generate distinct descriptions for the target categories when their names are abstract and technical, e.g. in NEU-DET dataset; for further discussion please refer to the Appendix.
4.4 Results on Standard FSOD Datasets
To further evaluate the generalization of our method, we test our model on a common FSOD benchmark, i.e. few-shot PASCAL VOC, in Tab. 2. Following TFA [56], as described in Sec. 4.2, there are three fixed groups of base-novel category splits, namely split 1, 2, and 3 to avoid sampling bias. For each data split, the $n$-shot ($n\in\{1,2,3,5,10\}$) sampling for the novel categories are also pre-defined. The model is evaluated on these three respective dataset splits. In addition, as described in Sec. 3.1, the base categories are also trained during fine-tuning to avoid forgetting, but have different sampling strategies: (i) in works such as TFA [56], the base categories also follow the restricted $n$-shot sampling (i.e. $n$ instances instead of $n$ images for each categories) as the novel categories; (ii) while in works such as Meta-DETR [60], the base categories are not few-shot. Both strategies are acceptable for the FSOD task; we adapt to the later one. As shown in Tab. 2, our method outperforms previous works.

### v2

@@ Abstract
Cross-modal feature extraction and integration have led to steady performance improvements in few-shot learning tasks. However, existing multi-modal object detection (MM-OD) methods degrade when facing significant domain shift and are sample insufficient. We hypothesize that rich text information could more effectively help the model to build a knowledge relationship between the vision instance and its language description and can help mitigate domain shift. Specifically, we study the Cross-Domain few-shot generalization of MM-OD (CDMM-FSOD) and propose a meta-learning based multi-modal few-shot object detection method that utilizes rich text semantic information as an auxiliary modality to achieve domain adaptation. Our proposed novel neural network contains (i) a multi-modal feature aggregation module that aligns the vision and language support feature embeddings and (ii) a rich text semantic rectify module that utilizes bidirectional text feature generation to reinforce multi-modal feature alignment and thus to enhance the model’s language understanding capability. We evaluate our model on common standard cross-domain object detection datasets and demonstrate that our approach considerably outperforms existing FSOD methods. Our implementation is publicly available: https://github.com/zshanggu/CDMM
@@ 1 Introduction
In real-world industrial application, the majority of current deep learning-based object detection approaches encounter failures when detecting product defects [40], primarily due to the scarcity of relevant training data and the substantial domain gap between pre-trained datasets from scenarios common in daily life versus the specific domain of product defect data. However, human experts are often equipped with comprehensive training manuals that provide detailed textual instructions and guidance, enabling them to discern object appearances and accurately annotate images. Consequently, in this paper, we propose a methodology that mimics this process, leveraging multi-modal rich textual information to augment the object detection task in scenarios where training data is scarce and out-of-domain.
@@
Figure 1: Different FSOD tasks. The classic FSOD task (top) involves only visual information. MM-FSOD (middle) introduces a language modality to provide extra information to improve FSOD performance. In contrast to these, our proposed CDMM-FSOD task (bottom) is tailored for cross-domain scenarios and extends MM-FSOD to use richer text. Above, we show a cross-domain example where a model might train on images and text of common data (such as birds) and needs to generalize to detection of less common, substantially different data (such as patch defects).
FSOD. We study the general few-shot object detection (FSOD) task. The goal is to train a model that can generalize to detecting objects for which we only have a few labeled examples [47, 39]. A typical FSOD model is first pre-trained on a number of base classes for which we have abundant annotated data and then the model is fine-tuned on a set of novel classes with few annotated samples. In the classic setting, the base and the novel classes are generally within the same feature domain. Multi-modal learning is a new approach for FSOD, and utilizes extra text information to enhance visual feature representations and provides potential zero-shot inference ability (see  Fig. 1). These methods encompass various techniques, including, visual question-answering  [16, 3, 52], robotics [2, 25], continual learning  [42, 4], and medical image analysis  [22, 31].
Cross-domain problem in FSOD. Despite the success of multi-modal object detection (MM-OD) methods on few-shot learning tasks, the source and the target data may have a substantial domain gap in many practical detection problems. Thus, the performance of MM-OD methods degrade in such cross-domain, data scarce scenarios. As shown in  Fig. 2, when we train and test object detection methods on data where the source and target domains differ, most MM-OD methods degrade significantly. Therefore, our goal is to address the cross-domain multi-modal few-shot object detection (CDMM-FSOD) task, aiming to bridge the domain gap while transferring information from the source domain to the target domain to address this challenge.
Rich text. We hypothesize that using rich text within a sentence can be more effective than relying solely on simple semantics, especially when images in the target domain require technical language to describe. However, for neural network models, how to utilize such rich text information for detection tasks remains unclear. We aim to investigate whether and how rich text can address such cross-domain data gaps. In Fig. 1, we illustrate the difference of our proposed CDMM-FSOD compared to classic FSOD and MM-FSOD. Our rich text contains detailed technical descriptions of the training data, and our novel data (the patch defect in this example) has prominent domain shift. Given that this scenario is prevalent in industry, we believe our approach offers a way to tackle such practical challenges.
@@
Figure 2: Performance results on 10-shot object detection on multiple cross-domain, few-shot datasets. We observe substantial cross-domain degradation for existing MM-OD Next-chat [50], GOAT [46], and ViLD [13], as well as for a single-modal detection method, Meta-DETR [51]. In contrast, our proposed method has stronger performance on out-of-domain data.
@@
Our method. We choose a meta-learning version of Detection Transformer (DETR) [5], i.e. Meta-DETR [51] as our baseline because of its Transformer-based structure and excellent modality adaption capacity. Unlike the classic multi-modal setting, we define a rich text description for each training category which may contain technical terminology. These sentences function as the linguistic support set for our meta-learning process, analogous to the role played by the image support set in terms of functionality. We design a meta-learning multi-modal aggregated feature module to fuse the vision and the language embedding and map them to a class-agnostic feature space as the meta support feature. Furthermore, to ensure the model obtains the detailed knowledge from the rich semantics, we propose a rich semantic rectify module to align the generated language embedding with the ground truth language embedding. In our experiments, we demonstrate consider performance improvement by using rich text. We further discuss and investigate the effect of text length on performance.
Our specific contributions include:
@@
  • We propose a rich semantic rectify module that learns to align the generated language embedding and the ground truth language embedding, enabling the model to better understand the text for knowledge transfer.
  • Our experiments on cross-domain detection data demonstrate that our approach outperforms existing methods, suggesting that multi-modal learning with rich text is effective for detection of few-shot, out-of-domain data.
@@ 2.1 Few-shot Object Detection
The FSOD task is to train a model to accurately detect novel categories where only a few labeled samples are available. The dominant training paradigms of FSOD include methods based on fine-tuning and meta-learning. Fine-tuning methods often first pre-train a model on base data with sufficient training samples for each category, and then directly fine-tune the model on novel categories with few samples. Representative works include TFA [47], FSCE [43], FSRC [38], DeFRCN [32], etc. In meta-learning methods, i.e. learning to learn, instead of aiming at specific classes, a model is trained to learn a metric ability through tasks (i.e. episodes), so that it can utilize a Siamese network [23] to achieve accurate classification by measuring the similarity between support and query features; thus, it is expected to learn class-agnostic knowledge [6]. Each training episode in meta-learning methods builds a support set and query set; an $n$-way $k$-shot support set consists of $n$ categories with $k$ training instances for each. Such work includes Meta-FRCN [15], Meta-DETR [51], etc.
To avoid catastrophic forgetting [41, 10, 36] of the base categories, researchers tune FSOD models on samples of both base and novel classes. However, while fine-tuning methods directly transfer the model on novel data in a class-specific manner, meta-learning methods tend to result in class-agnostic models, because they aim at learning a metric scheme between query features and support features. TFA [47] demonstrates that fine-tuning methods have potential to surpass meta-learning methods in detection performance. However, FCT [17] shows that meta-learning is still a powerful solution for FSOD, and that such methods are better at rapidly adapting the model to new tasks, due to their task-level feature representation.
The traditional backbone networks for FSOD include Faster R-CNN [35], YOLO [34], and ViT [8]. More recently, Detection Transformer (DETR) [5, 37], is gaining increasing popularity for FSOD due to its competitive performance compared to classical frameworks; furthermore, its transformer-based architecture facilitates integration with NLP tasks. In our work, we thus follow the meta-learning paradigm provided by Meta-DETR [51].
@@ 3.2 Rich Text
We propose to manually define a fixed, rich text description for each category $c\in C_{B}\cup C_{N}$ to form the text $\ell$ for the training data for improved information content. We denote the set of text descriptions as $W^{C}=\{w^{1},w^{2},\ldots,w^{C}\}$, where $C=|C_{B}\cup C_{N}|$ is the number of categories. Each text description contains multiple tokens with variable length of $M$, noted as $w^{i}=\{w^{i}_{1},w^{i}_{2},\ldots,w^{i}_{M}\}$, where the first token $w^{i}_{1}=\langle s\rangle$ is the start of the sentence, and $w^{i}_{M}=\langle/s\rangle$ is the end of the sentence. To form the set $W^{C}$, we use two strategies: (i) we refer to the Wikipedia content of each category and define each category from the aspects of color, shape, attribute, material, etc. (ii) besides the aspects from the first strategy, we extend the sentence by adding the depiction of the image context while keep other text unchanged. For example, the original sentence “Motorbikes have two wheels, a motor, and a sturdy frame” will be extended as “Motorbikes have two wheels, a motor, and a sturdy frame, usually ride by a person.” In the revised sentence, we introduce a common visual relationship between two categories: motorbikes and persons.
Figure 3: The overall structure of our model. We indicate our proposed multi-modal feature aggregation module and rich text rectify module with the red blocks; see Fig. 4 for more details about their structure. The multi-modal feature aggregation module is responsible for the cross-modality feature embedding mix. The rich text rectify module reinforces the model’s cross-modality understanding. We design this end-to-end model that takes a set of support and query images, as well as a group of rich category text as input for training, and outputs the object detection results of the query images.
@@ 3.3 Proposed Architecture
We use Meta-DETR [51] as our baseline network, which is built upon deformable DETR [55]. DETR is a transformer-based [45] object detection model. It has a ResNet [18] backbone to obtain the high-level image features; and then an encoder is applied to get the encoder memory embedding; followed by a decoder that could convert the memory embedding back to the object proposal. For the backbone part, we follow Meta-DETR [51] and use a weight-shared self attention module to build the meta-learning scheme for the query and support images. The classic DETR network includes an encoder and a decoder module, both consists of six multi-head attention layers. Meta-DETR build the meta-learning module only at the first multi-head attention layer of the encoder module.
For the meta-learning module, the query and support images will be firstly sent to a shared ResNet feature extractor to get the basic image features. Then the query and support features will be passed through a multi-head attention module so that they are in the same feature space. The support sample feature will then serve as the $K$ and $V$ matrices while the query image feature is $Q$ of a single-head self-attention module. This single-head self-attention module is responsible for the class-specific feature matching among the support and query feature. Meanwhile, in order to align with the class-agnostic learning task setting of meta-learning, this single-head self-attention module will be applied on a parallel branch to map the class-specific ‘support’-‘query’ relationship to a class-agnostic ‘task embedding’-‘query’ relationship. Superficially, Meta-DETR designs an independent trainable class-agnostic feature prototype as $V^{\prime}$ for the parallel single-head self-attention branch. And the support and query sample feature will still serve as $K^{\prime}$ and $Q^{\prime}$ respectively. In this way, the single-head self-attention module will achieve an important feature mapping so that the original class-specific support sample feature could be converted to class-agnostic prototype feature. During the test time, the model will only use the class-agnostic prototype as support features. And then the aggregated class-agnostic prototype will be sent to the input of DETR to generate the final detection output.
@@
Figure 4: Details of our meta-learning multi-modal aggregation module (upper region) and the rich semantic rectify module (lower region). We use different colors for different feature branches. The rich semantic rectify module is only used during training, and not at test time.
@@
The overall architecture of our proposed method is shown in Fig. 3. We add the rich text label as the auxiliary support feature. Then we aggregate the image and language support features to train the single-head self-attention meta-learning module. This module is the meta-learning multi-modal feature aggregation module (MM Aggre.). Besides, in order to ensure that the model properly learn the knowledge from the rich text information, we adopt a rich text semantic rectify module (Rich Text Rect.) to bidirectionally generate language feature, and then enforce it to align with the ground-truth language feature.
@@
[…]
3.4 Multi-modal Feature Aggregation
Multi-modal feature aggregation is the core module to fuse the multi-modal information, which is capable of aggregating multiple support categories. We visualize its architecture in Fig. 4. We use the existing CLIP pre-trained model to extract the primary image feature of the support/query set and the primary semantic feature of the rich text. Then the support image feature will be processed by a RoIAlign layer and an average pooling layer to get the instance-level category prototypes. This is a common operation in object detection because the image feature could not represent the instance feature; there usually exists multiple foreground objects in one image. Simultaneously, similar to prior object detection works, the background, as an extra hidden label, is also initialized as one of the category prototypes. These features are then mapped to the same feature space through a shared multi-head self-attention module (i.e. the first DETR encoder layer); our ablation study in Sec. 4.5 verifies the importance of sharing weights. The rich text must be provided at inference time.
$A=\text{softmax}\left(\frac{QS^{T}}{\sqrt{d}}\right)$ (1)
$Q_{1}=A\sigma(S)\odot Q,$ (2)
where $\odot$ represents the element-wise Hadamard product. The output $Q_{1}$ is a refined query feature map that keeps only the ‘foreground’ features.
For the ‘task embedding’-‘query’ encoding matching, a set of class-agnostic task prototype ($T\in\mathbb{R}^{(C+1)\times d}$) is defined to replace the support class prototype. In other words, the class-specific support class prototype will be mapped to the pre-defined task prototype, so that the query feature will be matched with the task prototypes for the prediction process. The task prototype $T$ would serve as the $V^{\prime}$ matrix for the same single head self-attention module as the ‘support’-‘query’ feature mapping. The initialization of $T$ is similar to sinusoidal position embedding in transformer. The corresponding output would be:
$Q_{2}=AT$ (3)
3.5 Rich Text Semantic Rectify
The rich text semantic rectify module is depicted in  Fig. 4. Our idea is to use a transformer encoder structure  [1], noted by $f_{\theta}$ to generate the category-level language sequence features of the fused support-query samples, and then the generated language feature is expected to be aligned with its ground truth feature. Additionally, this language generation process is bidirectional, which means we generate the sentence in both forward and backward direction, so that we ensure the model has robust image-text understanding ability and encodes semantic knowledge. The Rich Semantic Rectify module is discarded at inference time.
For a given category $c$, its rich text tokens undergo refinement through the multi-head self-attention module that is shared with the vision branch to get its semantic prototype $l^{c}$, as shown in  Fig. 4. Additionally, the support and query features are averaged to create a composite feature $p$, which serves as the $K$ and $V$ matrices for a multi-head self-attention module $A_{3}$, as shown in  Fig. 4. In $A_{3}$, the refined text tokens act as the $Q$ matrix. Subsequently, it generates the predicted text sequence $\hat{W}$ of length $M$. This entire process is duplicated for both forward and backward prediction components, i.e. the model will predict $\hat{W}$ from both the left-to-right and the right-to-left directions so that it could understand the rich text through its bidirectional context. The loss function for this purpose would be:
$\begin{split}\mathcal{L}_{rect}=\frac{1}{2}\left[\sum_{i=1}^{C}\sum_{j=2}^{M}-\mathrm{log}f_{\theta}(\hat{w}^{i}_{j}-w^{i}_{j}|p)+\right.\\ \left.\sum_{i=1}^{C}\sum_{j=1}^{M-1}-\mathrm{log}f_{\theta}(\hat{w}^{i}_{j}-w^{i}_{j}|p)\right]\end{split}$ (4)
4 Experiments
4.1 Implementation Details
We conduct our experiments on hardware consisting of 8 3090 GPUs for parallel training and testing. We adopt hyper-parameter values from DETR [5] by setting the initial learning rate as 0.001, and using a batch size of 4 for 1-shot scenarios and 1 for 5-shot and 10-shot scenarios. The DETR encoder and decoder both have 6 self-attention layers (the first layer of the encoder is embedded in our meta-learning multi-modal aggregation module). Following the literature, we rely on Afham et al. [1] for the architecture of the image-text transformer decoder.